E16-15 (Weighted-Average Number of Shares) Newton Inc. uses a calendar year for financial reporting. The company is authorized to issue 9,000,000 shares of $10 par common stock. At no time has Newton issued any potentially dilutive securities. Listed below is a summary of Newton’s common stock activities. 6 Instructions (a) Compute the weighted-average number of common shares used in computing earnings per common share for 2013 on the 2014 comparative income statement. (b) Compute the weighted-average number of common shares used in computing earnings per common share for 2014 on the 2014 comparative income statement. (c) Compute the weighted-average number of common shares to be used in computing earnings per common share for 2014 on the 2015 comparative income statement. (d) Compute the weighted-average number of common shares to be used in computing earnings per common share for 2015 on the 2015 comparative income statement. 1. Number of common shares issued and outstanding at December 31, 2012 2,000,000 2. Shares issued as a result of a 10% stock dividend on September 30, 2013 200,000 3. Shares issued for cash on March 31, 2014 2,000,000 Number of common shares issued and outstanding at December 31, 2014 4,200,000 4. A 2-for-1 stock split of Newton’s common stock took place on March 31, 2015

Please demonstrate how to perform this problem
Answer #1

ANSWER

a)

Shares Periods outstanding stock dividend Aveage Shares outstanding

  2000000 1 january,2013 - 30 September,2013 (1+0.10) =1.1 2000000 * 1.1 * 9/12 =1650000

2200000 1 October,2013 - 31 December,2013 - 2200000 * 3 /12 = 550000

(after stock dividend) Total = 2200000

b)

Shares Periods outstanding Aveage Shares outstanding

2200000 1 January,2014 - 31 March,2014 2200000 * 3/12 = 550000

4200000 1 April,2014 - 31 December,2014 4200000 * 9/12 = 3150000   

        Total    3700000

c)

Shares Periods outstanding stock split Aveage Shares outstanding

   3700000 1 January,2014 - 31 December,2014 2 for 1 3700000 * 12/12 * 2 / 1 = 7400000

d )

Shares Periods outstanding stock split Aveage Shares outstanding

4200000 1 January,2015 - 31 March,2015 2 / 1 4200000 * 2 * 3/12 = 2100000

8400000 1 April,2015 - 31 December,2015 - 8400000 * 9 / 12 = 6300000 (After split up)     

  Total     8400000

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************